OpenStudy (anonymous):

Solve: x/(x-2) - 2/(2x-1 = 1/5 Any help is appreciated.

OpenStudy (saifoo.khan):

\[\frac{x}{x-2} - \frac{2}{2x-1} = \frac{1}{5}\]